OpenStudy (anonymous):

Nick is now seven times as old as Harry. In 6 years, Nick will be five times as old as Harry. How old will Harry be in 6 years?

OpenStudy (anonymous):

N=7*H N+6=5*(H+6) Substituting the first equation on the second: 7H+6=5H+30 7H-5H=30-6 2H=24 H=12 (This is Harry's current age) In 6 years, Harry will be 18

OpenStudy (anonymous):

let harrys present age = x years then nick is 7x years in 6 years harrys age = x+6 nicks age = 7x + 6 ATQ 7x + 6 = 5(x+6) => 7x + 6 = 5x + 30 => 7x - 5x = 30 - 6 => 2x = 24 => x = 24/2 = 12 so Harry's present age = 12 years After 6 years he will be 12+6 = 18 years